(7) Question 1 Convert 1.5 kg to mg: 1- 1500 mg 2- 2,900,000 mg 3- 1.5...

(7)

Question 1

Convert 1.5 kg to mg:

1-

1500 mg

2-

2,900,000 mg

3-

1.5 x 10^6 mg

4-

1.5 mg

Question 2

Convert 6.00 ft to cm:

1-

28.35 cm

2-

183 cm

3-

0.197 cm

4-

1.27 cm

Question 3

Convert 3.87 gallons to L, if there are 3.79 L in one gallon?

Question 4

Convert 4.57 in to nm, if there are 2.54 cm in one inch.
Enter the whole number (do not use scientific notation. Your answer will NOT follow the significant digit rules!)
Answer #1

1] 1Kg = 1000 gms

1gm = 1000 mg

1Kg = 10^6 mg

1.5 Kg = 1.5*10^6 mg

2] 1 ft =30.48 cm

6ft = 6*30.48 = 183 cm

3] 1 gallon = 3.79 L

3.87 gallons = 3.87*3.79 = 14.6673 L

4] 1 in = 2.54 cm

4.57 in = 11.607 cm

11.607 cm = 11.607*10^-2 metres

11.607*10^-2 *10^-9 / 10^-9 = 11.607*10^7 nm = 116070000